At the city museum, child admission is $5.40 and adult admission is $9.20. On Friday,172 tickets were sold for a total sales of $1221.40. How many child tickets were sold that day?
Answer by richwmiller(17219)

c+a=172,
5.4*c+9.2*a=1221.4
c=172-a
5.4*(172-a)+9.2*a=1221.4
928.8-5.4a+9.2*a=1221.4
3.8*a=292.6
a=77.0
c=172-a
c=95.0 a=77.0
check
95.0+77.0=172
5.4*c+9.2*a=1221.4
5.4*95.0+9.2*77.0=1221.4
513.0+708.4=1221.4
1221.4=1221.4
ok